OpenStudy (anonymous):

Cos(85)cos(35)-Sin(85)sin(35)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \cos(a+b)= \cos(a) \cos(b)-\sin(a) \sin(b) }\)

OpenStudy (solomonzelman):

your 'a' is `85`, and your 'b' is `35`

OpenStudy (anonymous):

I believe that I would use the trigonometric identity cos(u+or-v)=cosu(cosv)+or-sinu(sinv) @SolomonZelman

OpenStudy (anonymous):

That would come out to be cos(85+35)=cos(85)cos(35)-sin(85)sin(35) @SolomonZelman

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

And re-write \(\normalsize\color{blue}{ \cos(85+35) }\) as \(\normalsize\color{blue}{ \cos(120) }\)

OpenStudy (anonymous):

So that would then get me cos(120)=cos(85)cos(35)-sin(85)sin(35). Where do I go from there @SolomonZelman

OpenStudy (solomonzelman):

You can find the exact value of \(\normalsize\color{blue}{ \cos(120) }\) \(\normalsize\color{blue}{ \cos(120)=\cos(60+60)=\cos(60)\cos(60)-\sin(60)\sin(60) }\) \(\normalsize\color{blue}{ =\cos^2(60)-\sin^2(60) }\) 

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \cos^2(60)-\sin^2(60) = (\frac{1}{2})^2- (\frac{\sqrt{3}}{2})^2 =\frac{1}{4} -\frac{3}{4}=-\frac{2}{4}=-\frac{1}{2}}\)

OpenStudy (anonymous):

Is that the answer to the problem? @SolomonZelman

OpenStudy (solomonzelman):

yes as exact as I can get.

OpenStudy (anonymous):

Thank you very much @SolomonZelman

OpenStudy (anonymous):

How would the problem change if it was Cos(85)cos(35)+Sin(85)sin(35) instead of Cos(85)cos(35)-Sin(85)sin(35)? @SolomonZelman

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \cos(a±b)= \cos^2(a)∓\sin^2(a)}\)

OpenStudy (solomonzelman):

you tell me

OpenStudy (anonymous):

I don't know where to start @SolomonZelman Would the answer be positive instead of negative?

OpenStudy (solomonzelman):

it would be cos(85-35) instead of cos(85+35)

OpenStudy (solomonzelman):

you would get cos(50)

OpenStudy (anonymous):

Then what?

OpenStudy (solomonzelman):

then nothing, you can use your calculator to approximate it, but you can't really get an exact answer just like I did by cos(120)

OpenStudy (anonymous):

Is this using the sum and difference formula? @SolomonZelman

OpenStudy (solomonzelman):

What is using the sum or difference formula ?

OpenStudy (anonymous):

Nevermind, the sum and difference formula is just a trigonometric identity that my teacher wanted me to use when solving this problem. @SolomonZelman I appreciate your help, you have been very thorough and quick with your replies. Have a great day sir!